find the value of k so that the line contains (4, -3k) and (2k,5) has a slope of m = 5/2
Answer by stanbon(75887)

find the value of k so that the line contains (4, -3k) and (2k,5) has a slope of m = 5/2
-------
(5--3k)/(2k-4) = 5/2
------
Cross-multiply to get::
2(5+3k) = 5(2k-4)
10+6k = 10k-20
-----
4k = 30
k = 7.5
-------------
Cheers,
Stan H.